Given the following half cells: Fe2+/Fe0 and Pd/Pd2+, determine; a) what is the anode and what is the cathode, b) what is Eo, c)what is Go and k? Using the following concentrations, determine E value for your cell. Fe2+ [.25], Pd2+ [.05]. How long would you have to run this battery at 3.2 amps to lower the E by .05 volts?
Answer #1

We ean identfy cathode and moda fom valus Ancde厶か dec hode witんmore-ve Evalue. 2+ Fe S) Catho de yoen Reducher) b) 9 E = 1-3inc eases , When 1 .294 v = 1.355v-o-os9v 0654V 0.0706 V . 355-1.28 = log x 2>< 0.0706 V = = 2.39 02.39 = 245.47. F2t- PA 245